OpenStudy (anonymous):

answers An object is thrown off a cliff with a horizontal speed of 10m/sec. After 3 seconds the object hits the ground. Find the height of the cliff and the total horizontal distance traveled by the object.

OpenStudy (anonymous):

horizontal distance 30m and height will be 45m.

OpenStudy (anonymous):

Distance for falling is h = (1/2) (g) t^2 for g = 9.8 m/s^2, have h = 0.5 (9.8) (3)^2 = 44.1 m if use g=10, get 45 m, often accepted. horizontal distance = (10 m/s) (3 s) = 30 m